# The Signals of the Doomsday

### v3

@@ 6 Relativistic bubble-wall dynamics in a viscous medium and terminal velocity
If the bubble wall moves with the speed of light, then no signal can overtake it to warn us about the impending doom. However, a bubble interacts with its environment and can reach a terminal velocity that is less than the speed of light. Previously, this topic has been studied extensively in the literature, with works addressing bubble wall velocities, hydrodynamical constraints, and friction mechanisms in various contexts[42, 43, 44, 45, 46, 47, 48]. We begin by noting that a bubble wall propagating through matter or radiation (where some of it could be produced by the bubble itself) experiences friction due to particle interactions at its surface. Understanding this damping is crucial, as friction reduces the wall’s acceleration and sets its terminal velocity, thereby shaping the overall dynamics and determining the efficiency of associated particle production.
@@ 6.2 Terminal Velocity
$\Delta V\approx\eta\gamma.$
@@
$\gamma_{\text{term}}\approx\frac{\Delta V}{\eta}.$
@@
$v_{\text{term}}\approx 1-\frac{1}{2\gamma_{\text{term}}^{2}}.$
The friction coefficient $\eta$ can be calculated from microphysics. For a scalar field $\phi$ coupled to other particles (e.g., fermions) with mass $m\sim g\phi$, the friction can be estimated as:
$\eta\sim g^{2}\Gamma,$
where $g$ is the coupling constant, while $\Gamma$ is the interaction rate of particles with the wall. Suppose that the wall is moving in a hot dense plasma. In thermal equilibrium, $\Gamma\sim T^{4}$ (temperature of the plasma)[49], so
$\eta\sim g^{2}T^{4}.$
For the order of magnitude estimate, within the standard model, we can set $g\approx 0.1$. We can also generically set $\Delta V\sim T^{4}$. This gives
$v_{\text{term}}\approx 1-\frac{1}{2\gamma_{\text{term}}^{2}}=0.995c.$
This estimate may be acceptable in the early universe or in cases where the bubble is produced inside stars. However, in more complicated situations where the bubble is created in the late universe in vacuum, and encounters stars, planets, and diluted interstellar and intergalactic gas during its expansion, more detailed numerical simulations are warranted. After all, the bubble produces a large number of very high energy particles, so it is literally engulfed in its own plasma. Thus, precise calculations of the terminal velocity will be studied elsewhere, and we will instead consider several conservative values for illustration.
@@ 6.3 Proper acceleration of a spherical bubble wall in a viscous medium
which defines the characteristic time scale for the terminal velocity to be reached. As $R\to\infty$ the curvature term disappears and (6.14) reduces to
@@
The quantity $\tau_{\rm term}$ from (6.27) is important because particles will be produced predominantly up to this moment, when both the proper and coordinate accelerations become zero. In fact, particles might be produced after that moment via some other mechanism, but not due to this vacuum mismatch related to the Unruh effect, which requires a non-zero proper acceleration.
7 Modified Higgs particles production due to the friction
@@
$\sigma=5.8680\times 10^{-8}\;M_{p}^{3},\quad M=125.09\,\text{GeV}.$ (7.1)
@@
$A\equiv\frac{\Delta V}{\sigma},\qquad B\equiv\frac{\eta}{\sigma},\qquad A=1.000170415815\times 10^{-4}\;M_{p},$ (7.2)
@@
$\frac{dN_{\rm tot}}{d\tau}=N_{k=0}(\tau)\;4\pi R(\tau)^{2}\;\sinh y(\tau),\qquad N_{\rm tot}(0)=0.$ (7.8)
@@
$N_{\rm tot}(\tau)$ is the cumulative number produced up to time $\tau$ by integrating those increments beginning at nucleation. At the instant of nucleation $(\tau=0)$ no time has elapsed and no volume has been swept, so by definition the accumulated total is zero: $N_{\rm tot}(0)=0.$ $N_{k=0}(0)$ is, on the other hand, an instantaneous occupation per unit volume; we still need a finite swept volume (over a finite $d\tau$) to accumulate a non-zero total.

### v4

@@ 5 Relativistic bubble-wall dynamics in a viscous medium and terminal velocity
If the bubble wall moves with the speed of light, then no signal can overtake it to warn us about the impending doom. However, a bubble interacts with its environment and can reach a terminal velocity that is less than the speed of light. Previously, this topic has been studied extensively in the literature, with works addressing bubble wall velocities, hydrodynamical constraints, and friction mechanisms in various contexts[92, 91, 20, 77, 57, 18, 88, 79, 9, 19, 110, 94, 80, 100, 43]. We begin by noting that a bubble wall propagating through matter or radiation (where some of it could be produced by the bubble itself) experiences friction due to particle interactions at its surface. Understanding this damping is crucial, as friction reduces the wall’s acceleration and sets its terminal velocity, thereby shaping the overall dynamics and determining the efficiency of associated particle production.
@@ 5.1 Derivation of the equation of motion
Following the standard phenomenological treatment of relativistic bubble-wall damping, we model the frictional pressure entering Eq.(5.6) by the linear-drag form $P_{\rm fric}=\eta\,\gamma(v)\,v$ (see e.g. [81]). Ultrarelativistic microphysical calculations allow for different $\gamma$ -scalings of the thermal pressure, $P_{\rm th}\propto\gamma^{0,1,2},$ depending on the assumed kinematics and emission channels; we defer a detailed discussion of this scaling and its implications for our parametrization to Sec.7.
@@ 5.2 Terminal Velocity
$\Delta V\approx\eta\gamma.$ (5.10)
@@
$\gamma_{\text{term}}\approx\frac{\Delta V}{\eta}.$ (5.11)
@@
$v_{\text{term}}\approx 1-\frac{1}{2\gamma_{\text{term}}^{2}}.$ (5.12)
The friction coefficient $\eta$ can be calculated from microphysics. For a scalar field $\phi$ coupled to other particles (e.g., fermions) with mass $m\sim g\phi$, the friction coefficient produced by a relativistic plasma at temperature $T$ can be estimated on dimensional grounds as
$\eta\sim g^{2}T^{4},$ (5.13)
where $g$ is an effective coupling constant to the thermal bath.
In the late universe, we have more complicated situations where the bubble is created in the late universe in vacuum, and encounters stars, planets, and diluted interstellar and intergalactic gas during its expansion, more detailed numerical simulations are warranted. After all, the bubble produces a large number of very high energy particles, so it is literally engulfed in its own plasma. Thus, precise calculations of the terminal velocity will be studied elsewhere, and we will instead consider several conservative values for illustration.
@@ 5.3 Proper acceleration of a spherical bubble wall in a viscous medium
which defines the characteristic time scale for the terminal velocity to be reached. We justify the cutoff at $\tau_{\rm term}$ for the vacuum–mismatch channel and the choice $\tau_{\rm final}$ for the thermal channel in Appendix C. As $R\to\infty$ the curvature term disappears and (5.19) reduces to
@@
The quantity $\tau_{\rm term}$ from (5.32) is important because it marks the onset of the quasi–terminal regime. Vacuum–mismatch particle production is dominant during the interval $\tau\lesssim\tau_{\rm term}$, while for $\tau\gtrsim\tau_{\rm term}$ the proper acceleration has decayed on this timescale, so $N_{k=0}(\tau)$, and hence $\mathrm{d}N_{k=0}/\mathrm{d}\tau$, become exponentially small. The corresponding numerical bounds, which show that the late–time contribution to integrated particle yield beyond $3\tau_{\rm term}$ or $5\tau_{\rm term}$ is completely negligible, are presented in Appendix C. Particles may still be produced at later times via other mechanisms (as we will show below), but not through this Unruh–like vacuum–mismatch channel, which is controlled by a non–zero proper acceleration.
6 Particle production from vacuum mismatch limited by friction
@@
$\sigma=5.868\times 10^{-8}\;M_{p}^{3},\quad M=125.09\,\mathrm{GeV}.$ (6.1)
@@
$A\equiv\frac{\Delta V}{\sigma},\qquad B\equiv\frac{\eta}{\sigma},\qquad A=1.001\times 10^{-4}\;M_{p},$ (6.2)
@@
$\frac{dN_{k=0}}{d\tau}=N_{k=0}(\tau)\;4\pi R(\tau)^{2}\;\sinh y(\tau),\qquad N^{(\text{int})}_{k=0}(0)=0.$ (6.8)
@@
$N_{k=0}^{\rm(int)}$ is the cumulative number produced up to time $\tau$ by integrating those increments beginning at nucleation. At the instant of nucleation $(\tau=0)$ no time has elapsed and no volume has been swept, so by definition the accumulated total is zero. $N_{k=0}(0)$ is, on the other hand, an instantaneous occupation per unit volume; we still need a finite swept volume (over a finite $d\tau$) to accumulate a non-zero total.
Scenario $\eta\,[M_{p}^{4}]$ $\tau_{\rm term}\,[M_{p}^{-1}]$ $R_{\rm fin}\,[M_{p}^{-1}]$ $N_{k=0}^{\rm(int)}$
$\delta=10^{-8}$ $8.300\mathrm{e}{-16}$ $7.069\mathrm{e}{+07}$ $7.208\mathrm{e}{+07}$ $1.052\mathrm{e}{+08}$
$\delta=10^{-9}$ $624\mathrm{e}{-16} 2.2$$235\mathrm{e}{+08}$ $2.320\mathrm{e}{+08}$ $3.530\mathrm{e}{+09}$
$\delta=10^{-10}$ $8.300\mathrm{e}{-17}$ $7.069\mathrm{e}{+08}$ $7.147\mathrm{e}{+08}$ $1.136\mathrm{e}{+11}$
Table 1: For each ultra-relativistic deficit $\delta=1-v_{\rm term}$ we evolve the spherical wall with proper acceleration (6.4) and kinematics to the time $\tau_{\rm term}=\sigma/\eta$ where the terminal velocity is reached. The proper acceleration $\alpha(\tau)$ directly feeds the instantaneous occupancy (6.6); the total yield $N_{k=0}^{\rm(int)}$ follows from (6.8). We note that we took into account only the Higgs particles produced at rest, i.e. the $k=0$ mode. The number of particles produced in all momentum modes is higher than shown here.
The details of the numerical calculation have been described in Appendix B. The results show that as the radius grows, friction balances the drive, with the system approaching terminal velocity on the time-scale $\tau_{\rm term}$; and the integrated yield is dominated by the combination of the exponentially sensitive $N_{k=0}(\alpha)$ and the rapidly growing area factor $4\pi R^{2}$, which becomes substantial by the time $\tau\sim\tau_{\rm term}$ in the most ultra-relativistic cases. However, the production cut-off time $\tau_{\rm term}$ is very short. The number of particles shown in the Table 1 might be sufficient if the bubble is produced near us (after all a few detected particles at enormous energies will represent an intriguing signal), but their flux will be severely diluted when propagating over the cosmological distances. However, all the difference in the energy density between the false and true vacuum has to go somewhere. If friction practically quenched particle production due to the vacuum mismatch, most of that energy will go into thermal particle production because of intense heat absorbed by the environment.
In this section we intentionally worked with an idealised setup where the drag coefficient $\eta$ is held fixed in time. In other words, the friction ratio $B=\eta/\sigma$ does not respond to how hot or dense the shocked plasma becomes. The point of this toy model is to cleanly isolate the “vacuum–mismatch” contribution to particle production: an accelerating bubble wall imposes a time–dependent boundary condition on the Higgs field, and the non–adiabatic change in the mode frequencies $\omega_{\pm}$ produces particles even if we do not allow the medium to thermalise. In this sense, the zero–mode occupation number $N_{k=0}(\tau)$ should be interpreted as a minimal, non-thermal lower bound on the Higgs–channel particle yield that follows purely from the acceleration history of the wall.
7 Thermal particle production from frictional dissipation
In this section we will use some of the quantities derived before, but will repeat them anyway for clarity. As the bubble wall propagates through an ambient medium, microscopic interactions between the wall and the surrounding particles exert a frictional pressure
$P_{\rm fric}=\eta\,\gamma v.$ (7.1)
This friction acts against the vacuum–mismatch force that normally accelerates the bubble wall, thereby reducing the kinetic energy that the wall would otherwise acquire. However, the missing kinetic energy is not lost: it is deposited directly into a thin layer of shocked plasma immediately behind the advancing wall. Because the wall moves ultra-relativistically, this shocked layer thermalizes essentially instantaneously compared to the timescale on which the wall evolves. Our goal is to compute how much thermal energy and, consequently, how many thermal particles are produced by this dissipation mechanism.
The derivation presented here follows a fully self-consistent energy-conservation framework. It does not assume any specific microphysical interaction; instead, it relies solely on the difference between the trajectory of a frictionless bubble wall and that of a friction-limited wall. This difference encodes all dissipative effects in a completely model-independent way. Our treatment is consistent with the relativistic fluid analysis of Landau & Lifschitz and with the microscopic studies of bubble-wall friction in first-order phase transitions [92, 91, 18, 7, 57, 12]. Thermalization within the shocked layer is treated using standard equilibrium thermodynamics [73, 3].
Astrophysical environments are highly diverse (interstellar gas, stellar matter, self-generated plasmas). The microscopic scattering rates and dominant interaction channels are not known in general. To avoid tying our analysis to any specific model, we parametrize all dissipation through the terminal-velocity deficit
$\delta\equiv 1-v_{\rm term},$ (7.2)
which measures how close the terminal velocity is to the speed of light. The corresponding effective friction coefficient is then fixed by the force-balance condition (in the planar limit):
$\Delta V=\eta\,\gamma_{\rm term}v_{\rm term}.$ (7.3)
This relation ensures that, for any chosen $\delta$, the wall indeed saturates at $v_{\rm term}$ due to friction which we also showed previously. Since $\eta(\delta)$ encapsulates all microphysical effects, this framework is fully model-independent. We focus on three representative values which we have used throughout the paper as a benchmark,
$\delta=10^{-8},\qquad 10^{-9},\qquad 10^{-10},$
which capture the deeply ultra-relativistic regime.
Local energy conservation and the correct energy deficit
The energy stored in the wall per unit area is (as before)
$E_{\rm wall}=\sigma\,\gamma(t),\qquad\gamma(t)=\frac{1}{\sqrt{1-v^{2}}},$ (7.4)
where $\sigma$ is the surface tension and $\gamma$ is the Lorentz factor. This quantity we also used in section 5. Importantly, curvature affects the wall’s acceleration, but does not change the form of this energy density.
To quantify dissipation, we compare two cases:
(i) The frictionless wall, which accelerates as strongly as the vacuum pressure allows;
(ii) The physical wall, whose acceleration is reduced by friction.
The frictionless wall obeys
$\sigma\,\gamma_{0}^{3}\frac{dv_{0}}{dt}=\Delta V-\frac{2\sigma}{R},$ (7.5)
where $v_{0}$ (and thus $\gamma_{0}$) corresponds to the wall’s velocity without friction, which comes enormously close to the speed of light, but is not exactly equal to it. The term $2\sigma/R$ represents the counteracting curvature force, and the remainder provides the net acceleration.
As we have already shown in section 5, the friction-limited wall satisfies
$\sigma\,\gamma^{3}\frac{dv}{dt}=\Delta V-\frac{2\sigma}{R}-\eta\,\gamma v,$ (7.6)
where the last term represents the energy lost to the medium per unit time per unit area.
The key identity follows by differentiating $\gamma=(1-v^{2})^{-1/2}$:
$\frac{d\gamma}{dt}=\gamma^{3}v\,\frac{dv}{dt}.$ (7.7)
This identity ties changes in kinetic energy directly to changes in velocity. Applying it to the frictionless trajectory gives
$\frac{d\gamma_{0}}{dt}$ $=\gamma_{0}^{3}v_{0}\frac{dv_{0}}{dt}=\gamma_{0}^{3}v_{0}\left(\frac{\Delta V-2\sigma/R}{\sigma\gamma_{0}^{3}}\right)$ (7.8)
$\Rightarrow\frac{d\gamma_{0}}{dt}=\frac{v_{0}}{\sigma}\left(\Delta V-\frac{2\sigma}{R}\right).$ (7.9)
An analogous expression follows for the frictional trajectory:
$\frac{d\gamma}{dt}=\frac{v}{\sigma}\left(\Delta V-\frac{2\sigma}{R}-\eta\gamma v\right).$ (7.10)
We now define the central object of this analysis:
$E_{\rm max}(t)=\sigma\,\gamma_{0}(t),\qquad E_{\rm wall}(t)=\sigma\,\gamma(t),$ (7.11)
whose difference
$\Delta E(t)=\sigma\,[\gamma_{0}(t)-\gamma(t)]$ (7.12)
represents the energy that the wall should have gained in the absence of friction but did not. This energy deficit is the true measure of dissipation.
After differentiation we get the master equation:
$\frac{d}{dt}\Delta E(t)=(v_{0}-v)\left(\Delta V-\frac{2\sigma}{R}\right)+\eta\,\gamma v^{2}.$ (7.13)
Both terms on the right hand side vanish when $\eta\to 0$.
Thermalisation of the deficit
Once energy flows into the medium, it thermalises inside a comoving layer of fixed proper thickness $\ell$, so the total thermal energy is
$E_{\rm th}(t)=\rho_{\rm th}(t)\,A(t)\,\ell,\qquad A(t)=4\pi R^{2}(t).$ (7.14)
Because the thermal energy per unit area is exactly the accumulated deficit, we have
$\rho_{\rm th}(t)\,\ell=\Delta E(t).$ (7.15)
Differentiating,
$\frac{d\rho_{\rm th}}{dt}=\frac{1}{\ell}\frac{d}{dt}\Delta E(t),$ (7.16)
and inserting Eq. (7.13),
$\frac{d\rho_{\rm th}}{dt}=\frac{1}{\ell}\left[(v_{0}-v)\left(\Delta V-\frac{2\sigma}{R}\right)+\eta\,\gamma v^{2}\right].$ (7.17)
Switching to proper time via $d\tau=dt/\gamma$,
$\frac{d\rho_{\rm th}}{d\tau}=\frac{\gamma}{\ell}\left[(v_{0}-v)\left(\Delta V-\frac{2\sigma}{R}\right)+\eta\,\gamma v^{2}\right].$ (7.18)
It is convenient to parameterise the proper thickness of the shocked layer in terms of the same microscopic scale that controls the wall structure itself. In a Higgs-like scalar potential, the true–vacuum mass $\mu^{2}\sim V^{\prime\prime}(\phi_{\rm true})$ also sets the characteristic width of the bubble wall, $\delta_{\rm wall}\sim 1/\mu$, i.e. the distance over which the field interpolates between the false and true vacua. It is therefore natural to assume that the region in which the fluid is strongly disturbed and thermalised by the wall cannot be parametrically thinner than the wall, and is instead of the same order. We thus take the comoving thickness of the thermalised layer to scale as
$\ell\;\sim\;\frac{1}{\mu}\,$, (7.19)
where $\mu$ is the Higgs mass in the true vacuum (and also sets the width of the potential barrier).
The heating rate per unit volume is obtained by dividing the energy flux into the layer by this thickness. Using the energy–deficit law,
$\frac{d\rho_{\rm th}}{d\tau}=\frac{1}{\ell}\left[\gamma(v_{0}-v)\left(\Delta V-\frac{2\sigma}{R}\right)+\eta\,\gamma^{2}v^{2}\right],$ (7.20)
and substituting $\ell=1/\mu$ yields the working evolution equation
$\frac{d\rho_{\rm th}}{d\tau}=\mu\,\gamma(v_{0}-v)\left(\Delta V-\frac{2\sigma}{R}\right)+\mu\,\eta\,\gamma^{2}v^{2}.$ (7.21)
In this form the microscopic input is encoded in the scale $\mu$, which fixes the effective thickness of the shocked layer and is directly tied to the underlying Higgs potential, while the macroscopic drag coefficient $\eta(0)$ controls how efficiently the wall motion pumps energy into this layer.
The detailed value of $\ell$ is therefore not an independent free parameter: different $\mathcal{O}(1)$ variations in the proportionality between $\ell$ and $1/\mu$ would only rescale the overall normalisation of $\rho_{\rm th}$ and $N_{\rm th}$ by an $\mathcal{O}(1)$ factor, without affecting the qualitative behaviour of the heating or the relative importance of the two contributions in the brackets. The two terms now have clear physical meaning: (i) the lost–acceleration heating proportional to $(v_{0}-v)$, and (ii) the direct drag heating proportional to $\eta$. Both are essential, and both grow dramatically as the wall becomes ultra–relativistic.
7.1 Thermal particle production
Once the thermal energy density is known, the temperature follows from equilibrium thermodynamics:
$T(\tau)=\left(\frac{30}{\pi^{2}g_{*}}\,\rho_{\rm th}(\tau)\right)^{1/4},$ (7.22)
with corresponding equilibrium number density
$n(\tau)=\frac{\zeta(3)}{\pi^{2}}\,g_{*}\,T(\tau)^{3}=\frac{\zeta(3)}{\pi^{2}}\,g_{*}^{1/4}\!\left(\frac{30}{\pi^{2}}\right)^{3/4}\!\left[\rho_{\rm th}(\tau)\right]^{3/4}.$ (7.23)
As the wall moves, it sweeps out the comoving volume
$dV=4\pi R^{2}(\tau)\,\sinh y(\tau)\,d\tau,$ (7.24)
so the instantaneous thermal production rate is
$\frac{dN_{\rm th}}{d\tau}=4\pi R(\tau)^{2}\sinh y(\tau)\,n(\tau)=4\pi R^{2}(\tau)\sinh y(\tau)\;\frac{\zeta(3)}{\pi^{2}}g_{*}^{1/4}\left(\frac{30}{\pi^{2}}\right)^{3/4}\left[\rho_{\rm th}(\tau)\right]^{3/4}.$ (7.25)
Therefore thermal production depends entirely on the integrated heating history,
$\rho_{\rm th}(\tau)=\int_{0}^{\tau}\!d\tau^{\prime}\;\mu\left[\gamma(\tau^{\prime})\,\big(v_{0}(\tau^{\prime})-v(\tau^{\prime})\big)\left(\Delta V-\frac{2\sigma}{R(\tau^{\prime})}\right)+\eta(\tau^{\prime})\,\gamma(\tau^{\prime})^{2}v(\tau^{\prime})^{2}\right].$ (7.26)
as dictated by the evolution of the energy deficit $\Delta E$. The numerical procedure associated with the thermal production has been discussed in the next subsection but for the details of the general numerical procedure please refer to Appendix B